OpenStudy (anonymous):

4- a-4/a+3 = a^2-2/a+3

OpenStudy (saifoo.khan):

\[4-\frac{a-4}{a+3}=\frac{a^2-2}{a+3}\]

OpenStudy (saifoo.khan):

this/

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):